why defferential equation "x*x' = 1" is non-linear ?
Answer by uma(370)


QUESTION:
why equation "x*x' = 1" is non-linear ?
------------------------------------------------------------------------
SOLUTION:

A linear equation is an equation involving only the sum of constants or products of constants and the first power of a variable.
Such an equation is equivalent to equating a first-degree polynomial to zero.

These equations are called "linear" because they represent straight lines in Cartesian coordinates.
A common form of a linear equation in two variables is y = mx + b, (e.g. y = 3x + 5).

Equations involving terms such as x2, y1/3, and xy are "non-linear".
-------------------------------------------------------------------------------
ANSWER TO YOUR QUESTION:

X*X = 1 IA NOT A LINEAR EQUATION.
BECAUSE X*X MEANS X^2, THAT IS THE POWER OF THE TERM IS 2.
SO THIS IS NOT A LINEAR RQUATION.

BYE.
PRASEENA.